# SignBot: Learning Human-to-Humanoid Sign Language Interaction

### v1

@@ 4.3 SignBot’s Cerebral: Sign Language Reasoning
Within our SignBot humanoid agent, the cerebral system controls high-level reasoning skills, enabling it to respond sign language and generate appropriate responses. An ideal embodied humanoid robot needs to have both cerebellar and cerebral capabilities, enabling it to interact effectively with its surrounding environment and other agents. The third stage of Figure 1 illustrates the cerebral system of the SignBot humanoid robot. This system processes the user’s input and generates an appropriate sign language response, which serves as the imitation goal for the control policy ($\mathbf{o_{t}^{c}}$ in Section 4.2).
@@
Implementation. SignBot utilizes a camera to observe the motions of sign language users and then stores this as a video input into its cerebral system. For communicating with sign language users in real-time, SignBot’s cerebral system is implemented by three models: a sign language translator understanding sign language contents, a sign language responser interpreting semantics and generating responses, and a sign language generator converting texts to the SMPLX format with Transformer-based models. We introduce these two models in the following:
@@
Sign Language Translator is implemented with a LLM [28]. In the pre-training stage, it extracts the sign language features from sign language videos and images, aligns them with the dimensions of the language model, and then inputs them into the model. In the fine-tuning stage, we utilize the translation text of the sign language dataset [44; 45; 46] to construct the supervised label for fine-tuning our translation model.
@@
Sign Language Responser is implemented using the DeepSeek-R1 API [29] due to its ability to seamlessly comprehend semantic information and facilitate multi-turn conversations. To simulate natural conversations with sign language users, we design a suitable prompt template. When generating a response, the system integrates the text-format sign language into the prompt and produces a contextually appropriate reply. A prompt example can be seen in Appendix D.1.
@@
Sign Language Generator autoregressively generates 3D sign motions from text input using a pretrained multilingual LM [47]. Firstly, we design a decoupled VQVAE tokenizer to map continuous sign motions into discrete tokens over upper body (UB), left hand (LH), and right hand (RH) movements [48; 49]. Given a $K$-frame sign sequence, we decompose it into three part-wise motion sequences based on the SMPL-X format: $\mathbf{C}^{u}\in\mathbb{R}^{K}$, where $u\in\{\text{UB},\text{LH},\text{RH}\}$. For each body part, we train a separate VQ-VAE comprising an encoder that projects the sequence into a latent space $\mathbf{C}^{u}_{f_{enc}}=\{c^{u}_{f,k}\}_{k=1}^{K}\in\mathbb{R}^{K_{f_{enc}}\times C}$, a decoder for reconstruction, and a learnable codebook $\mathbf{Z}_{u}\in\mathbb{R}^{N^{u}_{Z}\times d}$, where $N^{u}_{Z}$ represents the number of codes and $d$ denotes the code dimension. Then, for each pose, we can derive a set of discrete tokens $\hat{\boldsymbol{z}}_{1,\dots,K}^{u}=[\hat{z}^{u}_{1},\dots,\hat{z}^{u}_{K}]$, which searches for the nearest neighbor from the codebook $Z_{u}$:
$\hat{z}^{u}_{k}=\arg\min_{z_{n}\in\mathbf{Z}_{u}}\|s^{u}_{f_{enc},k}-z_{n}\|^{2},\quad n\in[1,N^{u}_{Z}],\quad\forall{u\in\{\text{UB},\text{LH},\text{RH}\}}$ (3)
@@
Given a text description $\boldsymbol{x}$, the generator retrieves word-level signs based on the $\boldsymbol{x}$ from external dictionaries made by the decoupled tokenizer [3]. These word-level signs are represented with discrete tokens $\hat{\boldsymbol{z}}_{1,\dots,K}^{u}$. We feed these tokens and text sequence $\boldsymbol{x}$ into the LM encoder at the same time. During decoding, we adopt a multi-head decoding strategy [3]. We design three language modeling heads, implemented as fully connected layers, to predict motion tokens for each body part from $m$ simultaneously at each step. The decoding process can be formulated as:
$P_{Dec}(\boldsymbol{o}^{y}|\mathbf{h})$ $=\prod_{k=1}^{K}\prod_{u}P_{Dec}(o_{k}^{y^{u}}|\boldsymbol{o}_{<k}^{y,\mathbf{h}})\ =\prod_{k=1}^{K}P_{Dec}(o_{k}^{y^{\text{UB}}}|\boldsymbol{o}_{<k}^{y,\mathbf{h}})P_{Dec}(o_{k}^{y^{\text{LH}}}|\boldsymbol{o}_{<k}^{y,\mathbf{h}})P_{Dec}(o_{k}^{y^{\text{RH}}}|\boldsymbol{o}_{<k}^{y,\mathbf{h}})$ (4)
@@
where $(\boldsymbol{o}_{<k}^{y^{u}},\mathbf{h})$ is the simplification of $\boldsymbol{o}_{<k}^{y,\mathbf{h}}$, $\boldsymbol{o}^{y}=\{o_{1}^{y^{\text{UB}}},o_{1}^{y^{\text{LH}}},o_{1}^{y^{\text{RH}}},\dots,o_{K}^{y^{\text{UB}}},o_{K}^{y^{\text{LH}}},o_{K}^{y^{\text{RH}}}\}$ and $\mathbf{h}=f_{\text{Enc}}(\boldsymbol{x},\boldsymbol{\hat{z}}^{u}_{1,\dots,K})$ denotes the output of LM Encoder. Finally, the derived motion tokens are used to reconstruct sign motions. The training objective is to maximize the log-likelihood $\sum_{(\boldsymbol{x},\boldsymbol{o}^{y})\in\mathcal{D}}\log P_{\mathcal{G}}(\boldsymbol{o}^{y}\mid\boldsymbol{x})=\sum_{(\boldsymbol{x},\boldsymbol{o}^{y})\in\mathcal{D}}P_{Dec}(\boldsymbol{o}^{y}|f_{\text{Enc}}(\boldsymbol{x},\boldsymbol{\hat{z}}^{u}_{1,\dots,K}))$ over datasets $\mathcal{D}$.
The imitation goal produced by the generator corresponds to the sign language poses that the humanoid robot must perform. By using these poses ($\boldsymbol{o}_{t}^{y}$ in Section 4.2) to guide the control policy, users can interact effectively with SignBot in sign language.
@@
Sim2Real Deployment. In real-world environments, robot movement from positions from $t$ to $t+1$ is not instantaneous. To ensure smooth transitions, we employ the Ruckig algorithm [50] for online trajectory generation with third-order (jerk) constraints and complete kinematic targets. Ruckig computes time-optimal trajectories between arbitrary states, defined by position, velocity, and acceleration, while respecting velocity, acceleration, and jerk limits. To maintain smooth motion, we interpolate intermediate values between target positions. For safety during deployment, we avoid commanding joint angles near their physical limits, as slight discrepancies between simulation and reality may lead to motor power loss or low-voltage issues, even when using identical limit settings.
5 Environment
@@
Experiment Settings. To conduct a comprehensive evaluation, we quantify the performance of SignBot in simulated (IssacGym [51]) and realistic environments from the following perspectives: 1) Accuracy: How to effectively align sign language actions between robots and humans? 2) Generalization: How well does SignBot perform with diverse sign language datasets and different robots? 3) Naturalness: How effectively does SignBot imitate human-like sign language norms in the real-world interaction scenarios? We employ consistent evaluation metrics across all environments. These metrics include 1) Tracking Dof Pos Error tracks the errors of the demonstrations’ degree of freedom position. 2) Tracking Yaw tracks the error of the body yaw angle. 3) Tracking Linear Velocity tracks the error of the intrinsic linear velocity. 4) Tracking Roll&Pitch: tracks the errors of the body roll&pitch angle. 5) Cumulative Rewards: are calculated by all the weighted reward functions. We use the publicly available CSL-Daily Chinese Sign Language dataset [44] and the How2Sign American Sign Language dataset [46] as experimental training data. Our experiments are conducted across different embodiments. We utilize legged robots H1, wheeled robots W1, and Linker hand to verify our methods. Appendix A records details of our experimental setting.
@@
Comparison Methods. To demonstrate the effectiveness of SignBot, we compare with other baselines based on whole-body control or RL as follows: 1) SignBot (w/o Lower-Body Tracking follows SignBot, but allows the lower body to self-adapt and maintain balance. 2) Whole-Body Tracking + AMP like [19] uses an AMP reward [42; 52] to encourage the transitions of the policy to be similar to the motions of the sign language features. 3) Whole-Body Tracking like [20] learns the movement of the upper body and lower body simultaneously. Appendix B reports the detailed parameter settings.
5.1 Accuracy: Alignment of Sign Language between Human and Robot
Precise execution of sign language is vital for effective communication within the DHH community, as even minor inaccuracies can lead to misunderstandings, in contrast to boxing techniques, where approximate gesture replication suffices for demonstration purposes. We divide the CSL-Daily sign language dataset [44] into three difficulty levels according to the length of the sentences: simple (929 sentences), intermediate (4558 sentences), and difficult (1089 sentences). The levels of difficulty increase with the length of the evaluated sentence. We compare SignBot with previously mentioned baselines on the data across these three difficulty levels. Appendix B records the running details of this experiment.
@@ 5.1 Accuracy: Alignment of Sign Language between Human and Robot
Whole-Body Tracking + AMP $1.13\pm 0.01$ $0.11\pm 0.00$ $0.22\pm 0.00$ $0.04\pm 0.01$ $8.89\pm 0.32$
@@
Whole-Body Tracking + AMP $1.11\pm 0.02$ $0.10\pm 0.00$ $0.23\pm 0.01$ $0.04\pm 0.01$ $9.04\pm 0.24$
@@
Table 1: Tracking Performance: We compare model performance under three difficulty levels. The mean ± std results for each baseline are shown in the dataset under the three random seeds. Bolded results indicate the best performance. The rewards calculation denotes the cumulative returns over a trajectory, while the other metrics are calculated using the mean square error.
@@
Table 1 illustrates the training performance of each baseline under different difficulty levels. The results indicate that SignBot achieves the lowest error in metrics such as DoF position tracking and yaw angle tracking, while also achieving the highest reward value, demonstrating the effectiveness of the SignBot control policy. The improvement in tracking accuracy and the reduction in training difficulty are attributed to the provision of upper body DoF position data and the tracking of lower body bent standing position. SignBot excels in controlling yaw and roll&pitch angles, indicating strong stability of the robot base. We also test other baselines, such as tracking the upper and whole body DoF positions, but the performance of these is unsatisfactory. This is because sign language differs from other upper-body movements; it is flexible and varied, and the frequency of sign language actions in the dataset is relatively fast. In addition, we observe that the length of the sentences does not significantly affect SignBot’s performance. When the sentences are short, SignBot tends to have larger errors in tracking keypoints. This may be due to the shorter episode length, which causes the robot to complete and reset its random learning for the next action quickly. During this period, it needs to adjust the robot’s global pose frequently. We also provide a more intuitive case study to compare various baselines (See Appendix C). Please see more video examples.
5.2 Generalization: Imitation across diverse datasets and robots
@@
Generalizing SignBot to different embodiments. We retarget the human sign language data to align with the joints of humanoid robots to visualize the precision in imitating sign language poses.
@@
Figure 2: Sign Language Alignment between Human and Robots: We display the source video of human sign language in the first row, followed by the mesh from video processing, then the retargeted sign language demonstration on the SignBot, and the last row shows the result after training in Isaac Gym. In the last row, the red nodes represent the joints of the SignBot, while the green nodes represent the targetted demonstration nodes.
Notably, since the Unitree H1’s wrist has only one DoF, it can only rotate while maintaining a fixed orientation and cannot bend. To address this, we modify the robot by adding two additional DoFs to the wrist, which allows the wrist to move up/down and left/right, enabling more flexible sign language gestures. Figures 2 and 3 illustrate the alignment between human and robot sign language gestures after applying our SignBot method, demonstrated on both the legged Unitree H1 and the wheeled W1 robots. A notable observation is that SignBot maintains a high accuracy in imitating sign language across both robots. In particular, balancing is challenging for legged humanoid robots. To address this, we design a reward function that encourages both stability and accurate imitation of upper-body movements. By combining this reward formulation with our SignBot, the RL controller enables the robot (red joint) to track demonstration motions green joint).
Figure 3: Sign Language Alignment between Human and Wheeled Robot.
Generalizing SignBot to different Datasets.To demonstrate the generalization capability of SignBot, we evaluate SignBot with another English sign language dataset.
@@
SignBot $\boldsymbol{0.63\pm 0.02}$ $\boldsymbol{0.07\pm 0.02}$ $0.19\pm 0.03$ $0.07\pm 0.01$
@@
SignBot $\boldsymbol{0.57\pm 0.02}$ $\boldsymbol{0.11\pm 0.01}$ $0.17\pm 0.02$ $0.05\pm 0.00$
@@
SignBot $\boldsymbol{0.56\pm 0.01}$ $\boldsymbol{0.11\pm 0.02}$ $0.18\pm 0.02$ $\boldsymbol{0.05\pm 0.01}$
Table 2: Generalization Experiment.
Specifically, we preprocess a portion of the 2,286 data entries from the How2Sign dataset [46] and evaluate the robustness of the policy for different sign languages following Table 1. Table 2 illustrates SignBot’s performance on the How2Sign dataset, demonstrating that SignBot achieves a minimal tracking error while maintaining high rewards.
5.3 Naturalness: Sim-to-Real Human-Robot Interaction
To demonstrate the naturalness of SignBot, we demonstrate its performance in realistic environments. To deploy SignBot on the H1 robot, we must add a pair of wrists and an intermediary connection between the wrists and the arm. We find updating the hardware system challenging. To prevent hardware limitations, we conduct the real-world experiment on the Linker Hand and the W1 robot. The W1 robot and Linker Hand are controlled through the ROS system to drive the various joints during real deployment.
Figure 4 illustrates several examples of interactions between sign language users and a robot functioning as a supermarket cashier. In these scenarios, the robot initiates communication by asking customers about their intended purchases. As is typical in retail environments, the robot promotes the distinctive qualities of the products, such as the freshness of the fruits, to encourage customer engagement and stimulate sales. Subsequently, the robot addresses pricing inquiries by estimating realistic, everyday prices in its responses. Throughout these human-robot interactions, the robot exhibits a high degree of accuracy in executing critical sign language gestures, including pointing forward, giving a thumbs-up, and performing the "OK" sign. While occasional delays may occur, primarily for reasoning purposes or to ensure adherence to path planning constraints, the control policy ensures that each joint reaches its designated position along the planned trajectory. Due to space constraints, we are unable to present additional examples within the paper; however, further instances can be found in the supplementary video materials and Appendix D.

### v3

@@ IV-C SignBot’s Cerebral: Sign Language Reasoning
Within our SignBot humanoid agent, the cerebral system controls high-level reasoning skills, enabling it to respond to sign language and generate appropriate responses. An ideal embodied humanoid robot needs to have both cerebellar and cerebral capabilities, enabling it to interact effectively with its surrounding environment and other agents. The third stage of Figure 2 illustrates the cerebral system of the SignBot. This system processes the user’s input and generates an appropriate sign language response, which serves as the imitation goal for the control policy ($\mathbf{o_{t}^{c}}$ in Section IV-B).
@@
Implementation. SignBot utilizes a camera to observe the motions of sign language users and then stores this as a video input into its cerebral system. For communicating with sign language users in real-time, SignBot’s cerebral system is implemented by three models: a sign language translator understanding sign language contents, a sign language responder interpreting semantics and generating responses, and a sign language generator converting texts to the SMPLX format with Transformer-based models. We introduce these two models in the following:
@@
Sign Language Translator is implemented with a LLM [9]. In the pre-training stage, it extracts the sign language features from sign language videos and images, aligns them with the dimensions of the language model, and then inputs them into the model. In the fine-tuning stage, we utilize the translation text from the sign language dataset [16, 17] to construct the supervised labels for fine-tuning our translation model.
@@
Sign Language Responder is implemented using the DeepSeek-Chat API [10] due to its ability to seamlessly comprehend semantic information and facilitate multi-turn conversations. Other chat models can also be used here. To simulate natural conversations with sign language users, we design a suitable prompt template. When generating a response, the system integrates the text-format sign language into the prompt and produces a contextually appropriate reply. We ensure that DeepSeek is aware of the vocabulary and scope covered by the CSL dataset to prevent the output of dangerous or out-of-distribution (OOD) corpus.
@@
Sign Language Generator autoregressively generates sign motions from text input based on a multilingual LM [18]. First, we design a decoupled VQVAE tokenizer to map continuous sign motions to discrete tokens over upper body (UB), left hand (LH), and right hand (RH) movements. Given a $K$-frame sign sequence, we decompose it into three part-wise motion sequences based on the SMPL-X: $\mathbf{C}^{u}\in\mathbb{R}^{K}$, where $u\in\{\text{UB},\text{LH},\text{RH}\}$. For each body part, we train a separate VQ-VAE comprising an encoder that projects the sequence into a latent space $\mathbf{C}^{u}_{f_{enc}}=\{c^{u}_{f_{enc},k}\}_{k=1}^{K}\in\mathbb{R}^{K_{f_{enc}}\times C}$, a decoder for reconstruction, and a learnable codebook $\mathbf{Z}_{u}\in\mathbb{R}^{N^{u}_{Z}\times d}$, where $N^{u}_{Z}$ represents the number of codes and $d$ denotes the code dimension. Then, for each pose, we can derive a set of discrete tokens $\hat{\boldsymbol{z}}_{1,\dots,K}^{u}=[\hat{z}^{u}_{1},\dots,\hat{z}^{u}_{K}]$, which searches for the nearest neighbor from the codebook $Z_{u}$:
$\hat{z}^{u}_{k}=\arg\min_{z_{n}\in\mathbf{Z}_{u}}\|s^{u}_{f_{enc},k}-z_{n}\|^{2},\quad n\in[1,N^{u}_{Z}]$ (2)
@@
where $\forall{u\in\{\text{UB},\text{LH},\text{RH}\}}$. Given a text description $\boldsymbol{x}$, the generator retrieves word-level signs based on the $\boldsymbol{x}$ from external dictionaries made by the decoupled tokenizer. These word-level signs are represented with discrete tokens $\hat{\boldsymbol{z}}_{1,\dots,K}^{u}$. We feed these tokens and text sequence $\boldsymbol{x}$ into the LM encoder at the same time. During decoding, we adopt a multi-head decoding strategy. We design three language modeling heads, implemented as fully connected layers, to predict motion tokens for each body part from $m$ simultaneously at each step. The decoding process can be formulated as:
$P_{Dec}$ $(\boldsymbol{o}^{y}|\mathbf{h})=\prod_{k=1}^{K}\prod_{u}P_{Dec}(o_{k}^{y^{u}}|\boldsymbol{o}_{<k}^{y,\mathbf{h}})\$(3)
@@
where $(\boldsymbol{o}_{<k}^{y^{u}},\mathbf{h})$ is the simplification of $\boldsymbol{o}_{<k}^{y,\mathbf{h}}$, $\boldsymbol{o}^{y}=\{o_{1}^{y^{\text{UB}}},o_{1}^{y^{\text{LH}}},o_{1}^{y^{\text{RH}}},\dots,o_{K}^{y^{\text{UB}}},o_{K}^{y^{\text{LH}}},o_{K}^{y^{\text{RH}}}\}$ and $\mathbf{h}=f_{\text{Enc}}(\boldsymbol{x},\boldsymbol{\hat{z}}^{u}_{1,\dots,K})$ denotes the output of LM Encoder. Finally, the derived motion tokens are used to reconstruct sign motions.
@@
Sim-to-Real Deployment. In real-world environments, robot movement from positions from $t$ to $t+1$ is not instantaneous. To ensure smooth transitions, we employ the Ruckig algorithm [19] for online trajectory generation with third-order (jerk) constraints and complete kinematic targets. Ruckig computes time-optimal trajectories between arbitrary states, defined by position, velocity, and acceleration, while respecting velocity, acceleration, and jerk limits. To maintain smooth motion, we interpolate intermediate values between target positions. For safety during deployment, we avoid commanding joint angles near their physical limits, as slight discrepancies between simulation and reality may lead to motor power loss or low-voltage issues, even when using identical limit settings.
V Environment
@@
Experiment Settings. To conduct a comprehensive evaluation, we quantify the performance of SignBot in simulated (IssacGym [20]) and realistic environments from the following perspectives: 1) Accuracy: How to effectively align sign language actions between robots and humans? 2) Generalization: How well does SignBot perform with diverse sign language datasets and different robots? 3) Naturalness: How effectively does SignBot imitate human-like sign language norms in the real-world interaction scenarios? 4) Interactivity: What is the DHH community’s genuine evaluation of this work? We use the public CSL-Daily [16] and How2Sign (ASL) [17] Sign Language dataset. Experiments span different embodiments: H1 legged robot, W1 wheeled robot, and Linker hand.
Metrics. We adopt consistent metrics across settings, including errors in 1) DoF positions, body yaw, linear velocity, and roll/pitch. 2) Cumulative Rewards: are calculated by all the weighted reward functions. 3) BLEU-1 calculates the single-word overlap between the generated text and the reference. 4) BLEU-4 evaluates the match of four-word sequences (4-grams). 5) ROUGE assesses how well the generated output covers the essential content of the reference. 6) DTW-PA-JPE [13] evaluates sequence-level distances between the generated signs and ground truth.
@@
Comparison Methods. To demonstrate the effectiveness of SignBot, we compare with other baselines based on whole-body control or RL as follows: 1) SignBot (w/o Lower-Body Tracking follows SignBot, but allows the lower body to self-adapt and maintain balance. 2) Whole-Body Tracking + AMP uses an AMP reward [14] to encourage the transitions of the policy to be similar to the motions of the sign language features. 3) Whole-Body Tracking [5] learns the movement of the upper body and lower body simultaneously.
V-A Accuracy: Alignment of Sign Language between Human and Robot
Precise execution of sign language is vital for effective communication within the DHH community, as even minor inaccuracies can lead to misunderstandings, in contrast to boxing techniques, where approximate gesture replication suffices for demonstration purposes. To construct a high-fidelity dataset with precise 3D SMPLX annotations, we leverage state-of-the-art methods for 3D hand [21] and body reconstruction [22]. Specifically, for each 2D sign language video, we first utilize OSX [22] to obtain an initial body pose estimation. Recognizing that OSX often struggles with accurate arm and hand pose estimation, we implement a two-stage refinement process. For high-precision hand pose reconstruction, we employ WiLoR [21], a cutting-edge 3D reconstruction pipeline that can robustly detect and reconstruct even challenging hand configurations. The hand pose parameters and global orientation outputs from WiLoR are then directly integrated to replace the corresponding OSX-derived estimates, ensuring high fidelity.
As the datasets contain multiple demonstrators performing the same actions, we select the sign language motions of one demonstrator and segment them by difficulty based on the length of the label (longer labels correspond to motions of longer duration, for example, labels with 10 words, 10-20 words, and 20 words). We divide the CSL-Daily sign language dataset [16] into three difficulty levels: simple (929 sentences), intermediate (4558 sentences), and difficult (1089 sentences). We compare SignBot with the previously mentioned baselines on the data across these three difficulty levels. Appendix A.3 records the running details of this experiment.
@@ V-A Accuracy: Alignment of Sign Language between Human and Robot
Whole-Body Tracking + AMP $1.13\pm 0.01$ $0.11\pm 0.00$ $22\pm 0.00 0.0$$04\pm 0.01$ $8.89\pm 0.32$
@@
Whole-Body Tracking + AMP $1.11\pm 0.02$ $0.10\pm 0.00$ $23\pm 0.01 0.0$$04\pm 0.01$ $9.04\pm 0.24$
@@
TABLE I: Tracking Performance: We compare model performance under three difficulty levels. Bolded results indicate the best performance. The rewards calculation denotes the cumulative returns over a trajectory, while the other metrics are calculated using the mean square error.
@@
Notably, since the Unitree H1’s wrist has only one DoF, it can only rotate while maintaining a fixed orientation and cannot bend. To address this, we modify the robot by adding two additional DoFs to the wrist, which allows the wrist to move up/down and left/right, enabling more flexible sign language gestures. Table I illustrates the training performance of each baseline under different difficulty levels. The results indicate that SignBot achieves the lowest error in metrics such as DoF position tracking and yaw angle tracking, while also achieving the highest reward value, demonstrating the effectiveness of the SignBot control policy. The improvement in tracking accuracy and the reduction in training difficulty are attributed to the provision of upper body DoF position data and the tracking of lower body bent standing position. SignBot excels in controlling yaw and roll&pitch angles, indicating strong stability of the robot base. We also test other baselines, such as tracking the upper and whole body DoF positions, but the performance of these is unsatisfactory. This is because sign language differs from other upper-body movements; it is flexible and varied, and the frequency of sign language actions in the dataset is relatively fast. In addition, we observe that the length of the sentences does not significantly affect SignBot’s performance. When the sentences are short, SignBot tends to have larger errors in tracking keypoints. This may be due to the shorter episode length, which causes the robot to complete and reset its random learning for the next action quickly. During this period, it needs to adjust the robot’s global pose frequently.
V-B Generalization: Imitation across diverse datasets and robots
@@
Generalizing SignBot to different embodiments. We retarget the human sign language data to align with the joints of humanoid robots to visualize the precision in imitating sign language poses. Figure 3 illustrates the alignment between human and robot sign language gestures after applying our SignBot method, demonstrated on both the legged Unitree H1 and the wheeled W1 robots. A notable observation is that SignBot maintains a high accuracy in imitating sign language across both robots.
@@
Figure 3: Sign Language Alignment between Human and Robots: We display the source video of human sign language in the first row, followed by the mesh from video processing, and the last row shows the results of different robots. For the H1 robot, the red nodes represent the robot dof pos, while the green nodes represent the retargeted demonstration nodes.
Generalizing SignBot to different Datasets.To demonstrate the generalization capability of SignBot, we also evaluate SignBot with the How2Sign dataset. Specifically, we preprocess a portion of the 2,286 data entries and evaluate the robustness of the policy for different sign languages following Table I. Table II illustrates SignBot’s performance on the How2Sign dataset, demonstrating that SignBot achieves a minimal tracking error.
@@
SignBot $\boldsymbol{0.63\pm 0.02}$ $\boldsymbol{0.07\pm 0.02}$ $19\pm 0.03 0.0$$07\pm 0.01$
@@
SignBot $\boldsymbol{0.57\pm 0.02}$ $\boldsymbol{0.11\pm 0.01}$ $17\pm 0.02 0.0$$05\pm 0.00$
@@
[…]
TABLE II: Generalization Experiment.
V-C Naturalness: Sim-to-Real Human-Robot Interaction
To demonstrate the naturalness of SignBot, we choose the W1 robot to demonstrate the performance of the SignBot framework in realistic environments. The W1 robot and Linker Hand are controlled through the ROS system to drive the various joints during real deployment. Figure 4 illustrates several examples of interactions between a human and a robot functioning as a supermarket cashier. The individual shown in the figure is a proficient sign language user from our research team. In this scenario, the robot initiates communication by asking customers about their intended purchases. More examples can be found in the supplementary video material.